1. The time spent reading an article by a visitor to an online website has a mean of 345 seconds with a standard deviation of 180 seconds. Suppose 3,400 website visitors read an article posted on this website.

(a) Compute the probability that the average amount of time spent reading the article by all 3,400 visitors is between 340 and 350 seconds.

(b) Compute the probability that the total amount of time spent reading the article by all 3,400 visitors is at least 325 hours (careful, you need to convert hours to seconds).
Answer #1

= 345

= 180

a) if n = 3400

Now calculate p(340 < x < 350)

= p(x-/(/sqrt (n) < z< 0) + p(0 < z < x-/(/sqrt (n))

= p(340-345/(180/sqrt(3400) < z <0) + p(0<z<350-345/(180/sqrt(3400))

= p(-1.62 < z <0) + p(0 < z < 1.62)

= 0.4474 + 0.4474

=0. 8948

P(340 < x < 350) = 0.8948